hate:

A ball is dropped from a certain height. The function below represents the height f(n), in feet, to which the ball bounces at the nth bounce: f(n) = 9(0.7)n What does the number 9 in the function represent?

Vocaloid:

f(n) = 9(0.7)^n think about what happens when you plug in n = 0. f(0) = 9(0.6)^0 = 9 in general, this is true for any exponential function in the form f(x) = ab^x. a is always the value at x = 0. so thinking about your problem, what is the real-world significance of the value 9?